### v1

# On total transitivity of graphs

Abstract
Let $G=(V,E)$ be a graph where $V$ and $E$ are the vertex and edge sets, respectively. For two disjoint subsets $A$ and $B$ of $V$, we say $A$ dominates $B$ if every vertex of $B$ is adjacent to at least one vertex of $A$. A vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ is called a transitive partition of size $k$ if $V_{i}$ dominates $V_{j}$ for all $1\leq i<j\leq k$. In this article, we study a variation of transitive partition, namely total transitive partition. The total transitivity $Tr_{t}(G)$ is equal to the maximum order of a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ obtained by repeatedly removing a total dominating set from $G$, until no vertices remain. Thus, $V_{1}$ is a total dominating set of $G$, $V_{2}$ is a total dominating set of the graph $G_{1}=G-V_{1}$, and, in general, $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{i}$. A vertex partition of order $Tr_{t}(G)$ is called $Tr_{t}$-partition. The Maximum Total Transitivity Problem is to find a total transitive partition of a given graph with the maximum number of parts. First, we characterize split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Moreover, for the split graph $G$ and $11\leq p\leq\omega(G)-$ 1, we give some necessary conditions for $Tr_{t}(G)=p$. Furthermore, we show that the decision version of this problem is NP-complete for bipartite graphs. On the positive side, we design a polynomial-time algorithm to solve Maximum Total Transitivity Problem in trees.
Keywords. Total transitivity, Split graphs, NP-completeness, Polynomial-time algorithm, Trees.
1 Introduction
Partitioning a graph is one of the fundamental problems in graph theory. In the partitioning problem, the objective is to partition the vertex set (or edge set) into some parts with desired properties, such as independence, minimal edges across partite sets, etc. In literature, partitioning the vertex set into certain parts so that the partite sets follow particular domination relations among themselves has been studied [1, 6, 8, 9, 13, 17, 18, 24]. Let $G$ be a graph with $V(G)$ as its vertex set and $E(G)$ as its edge set. When the context is clear, $V$ and $E$ are used instead of $V(G)$ and $E(G)$. The neighbourhood of a vertex $v\in V$ in a graph $G=(V,E)$ is the set of all adjacent vertices of $v$ and is denoted by $N_{G}(v)$. The degree of a vertex $v$ in $G$, denoted as $\deg_{G}(v)$, is the number of edges incident to $v$. A vertex $v$ is said to dominate itself and all its neighbouring vertices. A dominating set of $G=(V,E)$ is a subset of vertices $D$ such that every vertex $x\in V\setminus D$ has a neighbour $y\in D$, that is, $x$ is dominated by some vertex $y$ of $D$.
There has been a lot of research on graph partitioning problems based on a domination relationship between the different sets. Cockayne and Hedetniemi introduced the concept of domatic partition of a graph $G=(V,E)$ in 1977, in which the vertex set is partitioned into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that each $V_{i}$ is a dominating set of $G$ [4]. The number representing the highest possible order of a domatic partition is called the domatic number of G, denoted by $d(G)$. Another similar type of partitioning problem is Grundy partition. Christen and Selkow introduced a Grundy partition of a graph $G=(V,E)$ in 1979 [2]. In the Grundy partitioning problem, the vertex set is partitioned into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that each $V_{i}$ is an independent set and for all $1\leq i<j\leq k$, $V_{i}$ dominates $V_{j}$. The maximum order of such a partition is called Grundy number of $G$, denoted by $\Gamma(G)$. In 2018, Hedetniemi and Hedetniemi [13] introduced a transitive partition as a generalization of the Grundy partition. A transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i<j\leq k$, $V_{i}$ dominates $V_{j}$. The maximum order of such a transitive partition is called transitivity of $G$ and is denoted by $Tr(G)$. Recently, in 2020, Haynes et al. generalized the idea of domatic partition as well as transitive partition and introduced the concept of upper domatic partition of a graph $G$, where the set of the vertex is divided into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for each $i,j$, with $1\leq i<j\leq k$, either $V_{i}$ dominates $V_{j}$ or $V_{j}$ dominates $V_{i}$ or both [9]. The maximum order of such an upper domatic partition is called upper domatic number of $G$, denoted by $D(G)$. All these problems—domatic number [1, 26, 27], Grundy number [5, 6, 14, 24, 25], transitivity [8, 13, 17, 18], upper domatic number [9, 23] have been extensively studied both from an algorithmic and structural point of view. A Grundy partition is a transitive partition with the additional restriction that each partite set must be independent. In a transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, we require the domination property in one direction, that is, $V_{i}$ dominates $V_{j}$ for $1\leq i<j\leq k$. However, in an upper domatic partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, for all $1\leq i<j\leq k$, either $V_{i}$ dominates $V_{j}$ or $V_{j}$ dominates $V_{i}$ or both. The definition of each vertex partitioning problem ensures the following inequalities for any graph $G$. For any graph $G$, $1\leq\Gamma(G)\leq Tr(G)\leq D(G)\leq n$, where $n$ is the number of vertices of $G$.
Motivation: Many variations of the classical domination problem have been studied in the literature [10, 11, 12]. One of the natural variations is total dominating set, where our goal is to find a set $D$ such that every vertex of $V$ is dominated by a vertex of $D$. The concept of total domination was first formally introduced in 1980 by Cockayne et al. [3] as part of their study on domination in graphs. In the classical domination problem, we are interested in monitoring the vertices of a network that are not in a dominating set by the vertices of the dominating set. Total dominating ensures that every node in a network is adjacent to a control node, making it useful for designing fault-tolerant systems where every node has a backup connection. This variation of the domination problem has been well studied in the literature from both a structural and algorithmic point of view. It has been shown that finding the minimum total dominating set is NP-complete for general graphs, bipartite graphs [22], and split graphs [15], whereas it is solvable in polynomial time for trees [16].
The domatic number of a graph was introduced by E. J. Cockayne and S. T. Hedetniemi [4]. Furthermore, many variants of this concept were introduced and studied by B. Zelinka [30]. One such variation is total domatic number, which was introduced by Cockayne et al. [3]. There are several works done on total domatic number [28, 29]. In 2018, Hedetniemi and Hedetniemi [13] introduced the notion of transitive partition. In 2024, S. Paul and K. Santra introduced three variations of transitivity, namely, $2$-transitivity [20], $d_{2}$-transitivity [19], and strong transitivity [21]. The $2$-transitivity, $d_{2}$-transitivity, and strong transitivity are based on $2$-domination, $d_{2}$-domination, and strong domination, respectively. In this article, we explore another natural variation of the transitivity problem, focusing on total domination instead of domination. We introduce such a variation of transitivity problem, namely total transitive partition. Total transitivity $Tr_{t}(G)$ equals the maximum order of a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ obtained by repeatedly removing a total dominating set from $G$, until no vertices remain. Thus, $V_{1}$ is a total dominating set of $G$, $V_{2}$ is a total dominating set of the graph $G_{1}=G-V_{1}$, and, in general, $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{i}$. A vertex partition of order $Tr_{t}(G)$ is called $Tr_{t}$-partition. Note that every total transitive partition is also a transitive partition. Therefore, for any graph $G$, $1\leq Tr_{t}(G)\leq Tr(G)\leq n$. Since the total dominating set exists only when $G$ does not have any isolate vertex, in this paper, we focus only on graphs with no isolates.
An equivalent definition of total transitivity can be given in terms of domination among partite sets and the set itself. For two disjoint subsets $A$ and $B$, we say $A$ dominates $B$ if every vertex of $B$ is adjacent to a vertex of $A$. Again, a subset $A$ dominates itself if every vertex of $A$ is adjacent to another vertex of $A$. A total transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i\leq j\leq k$, $V_{i}$ dominates $V_{j}$. The maximum order of such a total transitive partition is called the total transitivity of $G$ and is denoted by $Tr_{t}(G)$.
We define another variation of transitivity, namely modified-total transitivity. For two disjoint subsets $A$ and $B$, we say $A$ m-totally dominates $B$ if every vertex of $A\cup B$ is adjacent to a vertex of $A$. A modified-total transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i<j\leq k$, $V_{i}$ m-totally dominates $V_{j}$. The maximum order of such a modified-total transitive partition is called the modified-total transitivity of $G$ and is denoted by $Tr_{t}^{m}(G)$. Later, we will use this parameter to solve total transitivity in a tree. Clearly, if $\pi$ is a total transitive partition, it is also a modified-total transitive partition, and hence, for a graph $G$, $Tr_{t}(G)\leq Tr_{t}^{m}(G)$. Also, note that in a modified-total transitive $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, $V_{k}$ can contain a single vertex, but if $\pi$ is a total transitive partition, $|V_{k}|\geq 2$.
The Maximum Total Transitivity Problem and its corresponding decision version are defined as follows.
Maximum Total Transitivity Problem Instance: A graph $G=(V,E)$ Solution: A total transitive partition of $G$ with maximum size
Maximum Total Transitivity Decision Problem Instance: A graph $G=(V,E)$, integer $k$ Question: Does $G$ have a total transitive partition of order at least $k$?
In this paper, we study this variation of transitive partition from a structure and algorithmic point of view. The main contributions are summarized below:
  1. (1.)
    We characterize split graphs with total transitivity equal to $1$ and $\omega(G)-1$.
  2. (2.)
    Given a split graph $G$ and integer $p$, $11\leq p\leq\omega(G)-$ 1, we give some necessary conditions for $Tr_{t}(G)=p$.
  3. (3.)
    We show that the Maximum Total Transitivity Decision Problem is NP-complete for bipartite graphs.
  4. (4.)
    We show that the Maximum Total Transitivity Problem can be solved in polynomial time for trees.
The rest of the paper is organized as follows. Section 2 contains basic definitions and notations that are followed throughout the article. This section also discusses the properties of total transitivity of graphs. In section 3, we characterize split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Also, in this section, we give some necessary conditions for $Tr_{t}(G)=p$ for split graphs $G$ and $11\leq p\leq\omega(G)-$ 1. Section 4 shows that the Maximum Total Transitivity Decision Problem is NP-complete in bipartite graphs. In Section 5, we design a polynomial-time algorithm for solving Maximum Total Transitivity Problem in trees. Finally, Section 6 concludes the article with some open problems.
2 Preliminaries
2.1 Notation and definition
Let $G=(V,E)$ be a graph with $V$ and $E$ as its vertex and edge sets, respectively. A graph $H=(V^{\prime},E^{\prime})$ is said to be a subgraph of a graph $G=(V,E)$ if and only if $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. For a subset $S\subseteq V$, the induced subgraph on $S$ of $G$ is defined as the subgraph of $G$ whose vertex set is, $S$ and edge set consists of all of the edges in $E$ that have both endpoints in $S$, and it is denoted by $G[S]$. The complement of a graph $G=(V,E)$ is the graph $\overline{G}=(\overline{V},\overline{E})$, such that $\overline{V}=V$ and $\overline{E}=\{uv|uv\notin E\text{ and }u\neq v\}$. The open neighbourhood of a vertex $x\in V$ is the set of vertices $y$ adjacent to $x$, denoted by $N_{G}(x)$. The closed neighborhood of a vertex $x\in V$, denoted as $N_{G}[x]$, is defined by $N_{G}[x]=N_{G}(x)\cup\{x\}$. A subset $D\subseteq V$ is said to be a total dominating set of $G$, if for every vertex $x\in V$, there exists a vertex $y\in D$ such that $xy\in E$. Breadth first search (BFS) is a graph searching technique that systematically explores all the vertices of a connected graph. The process starts at a designated vertex and explores all its neighbours before moving on to the next level of neighbours. The ordering in which the vertices are visited is called BFS ordering. A vertex is said to be a leaf vertex if its degree is $1$, and the vertex that is adjacent to a leaf vertex is said to be the support vertex of the leaf vertex.
A subset of $S\subseteq V$ is said to be an independent set of $G$ if no two vertices in $S$ are adjacent. A subset of $K\subseteq V$ is said to be a clique of $G$ if every pair of vertices in $K$ is adjacent. The cardinality of a maximum size clique is called clique number of $G$, denoted by $\omega(G)$. A graph is called bipartite if its vertex set can be partitioned into two independent sets. A graph $G=(V,E)$ is a split graph if $V$ can be partitioned into an independent set $S$ and a clique $K$.
2.2 Basic properties of total transitivity
In this subsection, we present some properties of total transitivity that also motivate us to study this variation of transitivity. First, we show the following bound for total transitivity.
Proposition 1.
For any graph $G$, $Tr_{t}(G)\leq min\{\Delta(G),\left\lfloor\frac{n}{2}\right\rfloor\}$, where $\Delta(G)$ is the maximum degree of $G$.
Proof.
Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a $Tr_{t}$-transitive partition of $G$. Let $x\in V_{k}$. As $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{i}$, $deg(x)\geq k$ and $V_{i}\geq 2$ for all $i$. Therefore, $Tr_{t}(G)=k\leq deg(x)\leq\Delta(G)$ and $Tr_{t}(G)\leq\left\lfloor\frac{n}{2}\right\rfloor$. Hence, $Tr_{t}(G)\leq min\{\Delta(G),\left\lfloor\frac{n}{2}\right\rfloor\}$. ∎
Next, we find the total transitivity of complete graphs, paths, cycles and complete bipartite graphs.
Proposition 2.
For the complete graph $K_{n}$ of $n$ vertices $Tr_{t}(K_{n})=\left\lfloor\frac{n}{2}\right\rfloor$.
Proof.
The partition $\{V_{1},V_{2},\ldots,V_{\left\lfloor\frac{n}{2}\right\rfloor}\}$ such that each set contains at least two vertices is a total transitive partition of $K_{n}$ of size $\left\lfloor\frac{n}{2}\right\rfloor$. Moreover, from the Proposition 1, we have $Tr_{t}(G)\leq min\{\Delta(K_{n}),\left\lfloor\frac{n}{2}\right\rfloor\}=min\{n,\left\lfloor\frac{n}{2}\right\rfloor\}=\left\lfloor\frac{n}{2}\right\rfloor$. Therefore, $Tr_{t}(K_{n})=\left\lfloor\frac{n}{2}\right\rfloor$. ∎
Remark 1.
From the definition of a total transitive partition, it is clear that it is also a transitive partition of $G$. But the difference $Tr(G)-Tr_{t}(G)$ can be arbitrarily large. From the Proposition 2, we have $Tr_{t}(K_{n})=\frac{n}{2}$ for $n$ is an even integer, and from [13], we know that $Tr(K_{n})=n$, hence the difference $Tr(K_{n})-Tr_{t}(K_{n})=\frac{n}{2}$ is arbitrarily large for $n$.
Proposition 3.
Let $P_{n}$ be a path of $n$ vertices, and then the total transitivity of $P_{n}$ is given as follows:
$Tr_{t}(P_{n})=\begin{cases}1&n=2\\ 2&n\geq 3\end{cases}$
Proposition 4.
Let $C_{n}$ be a cycle of $n$ vertices, then the total transitivity of $C_{n}$ is $2$.
Proposition 5.
For a complete bipartite graph $K_{m,n}$ with $m+n$ vertices $Tr_{t}(K_{m,n})=min\{m,n\}$.
Proof.
Let the vertex set of $K_{m,n}$ be $V=X\cup Y$ such that $|X|=m,|Y|=n$ and $m\leq n$. Consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{m}\}$, where $V_{i}=\{x_{i},y_{i}\}$, for all $2\leq i\leq m$ and $V_{1}$ contains all the other vertices of $K_{m,n}$. Since $K_{m,n}$ is a complete bipartite graph, $\pi$ is a total transitive partition of size $m$ and $Tr_{t}(K_{m,n})\geq m$. Next, we show that $Tr_{t}(K_{m,n})$ cannot be more than $m$. If possible, let $Tr_{t}(K_{m,n})=k\geq m+1$ and $\pi^{\prime}=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of $K_{m,n}$ of size $k$. Let $v\in V_{k}$, then the degree of $v$ must be at least $k\geq m+1$. So, the vertex $v$ cannot be from $Y$. Moreover, as $V_{k}$ dominates $V_{k}$, there exists $u\in V_{k}$ such that $uv\in E$ and $u\in Y$. Since $u\in V_{k}$, the degree of $u$ must be at least $k\geq m+1$. We have a contradiction as $deg(u)=m<k$. Therefore, $Tr_{t}(K_{m,n})\leq m$. Hence, $Tr_{t}(K_{m,n})=min\{m,n\}$. ∎
Similarly, as transitivity, if we union any two sets of a total transitive partition of order $k$, we create a total transitive partition of order $k-1$. Hence, we have a similar proposition as transitivity.
Proposition 6.
If $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ is a total transitive partition of a graph $G$, then for any two sets $V_{i}$ and $V_{j}$ in $\pi$, where $i<j$, the partition $\pi=\{V_{1},V_{2},\ldots,V_{i-1},V_{i}\cup V_{j},V_{i+1},\ldots,V_{j-1},V_{j+1},\ldots,V_{k}\}$ is a total transitive partition.
Given this, we have the following interpolation result.
Proposition 7.
Let $G$ be a graph and $k$ be the order of a total transitive partition of $G$, then for every $j$, $1\leq j\leq k$, $G$ has a total transitive partition of order $j$.
Proposition 8.
Let $G$ be a connected graph and $TTr(G)=k$, $k\geq 2$. Then there exists a total transitive partition of $G$ of size $k$, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, such that $|V_{k}|=2$.
Now we show that if the total transitivity of a graph $G$ is constant, that is, independent of the number of vertices, and $\pi=\{V_{1},V_{2},\ldots,V_{k-1},V_{k}\}$ is a total transitive partition of $G$, then the number of vertices of each $V_{i}$, for $2\leq i\leq k$, does not depend on $n$, rather it is bounded by some exponential in $i$.
Lemma 9.
Let $G$ be a connected graph and $Tr_{t}(G)=k$, $k\geq 3$. Then there exists a total transitive partition of $G$ of size $k$, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, such that $|V_{k}|=2$ and $|V_{k-i}|\leq 2^{i+1}$, for all $i$, $21\leq i\leq k-$ 2.
Proof.
Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of $G$ of size $k$, where $k\geq 3$. Let $x\in V_{k}$, as $V_{k}$ dominates $V_{k}$, there exists $y\in V_{k}$ such that $xy\in E(G)$. So, $|V_{k}|\geq 2$. Now if $|V_{k}-\{x,y\}|\geq 1$, then we move all the vertex except two, say vertices $x,y$, from the set $V_{k}$ to the first set $V_{1}$. After this transformation, we have, $\pi^{\prime}=\{V_{1}^{\prime}=V_{1}\cup(V_{k}-\{x,y\}),V_{2},\ldots,V_{k-1},V_{k}^{\prime}=\{x,y\}\}$ which is a total transitive partition of $G$ of size $k$ and $|V_{k}|=2$. Therefore, $\pi^{\prime}$ is a total transitive of $G$ of size $k$, such that $|V_{k}|=2$. Consider the base case when $i=1$. Note that to dominate $V_{k}$ and $V_{k-1}$, we need at most $2\cdot|V_{k}|$ vertices in, $V_{k-1}$ and the remaining vertices can be moved to, $V_{1}$ as in the previous argument. Therefore, $|V_{k-1}|\leq 2\cdot|V_{k}|=2\cdot 2=4=2^{i+1}$. Assume the induction hypothesis and consider that we have a total transitive partition $\pi=\{V_{1},V_{2},\ldots\,V_{k}\}$, such that $|V_{k}|=2$ and $|V_{k-j}|\leq 2^{j+1}$, $11\leq j\leq i-$ 1. Now consider the set $V_{k-i}$. Similarly, to dominate the $i+1$ sets that are following it and including it, we need at most $2\cdot|V_{k-i+1}|$ vertices in $V_{k-i}$. If $V_{k-i}$ contains more than those many vertices, the remaining vertices can be moved to $V_{1}$, maintaining total transitivity of $G$. Hence, $V_{k-i}\leq 2\cdot|V_{k-i+1}|=2\cdot 2^{i}=2^{i+1}$. ∎
When the number of vertices in each set is bounded by a constant, then the number of possible subsets that can be $V_{1}$ is polynomial in $n$, where $n$ is the number of vertices of the graph. The same is true for every set $V_{2},V_{3},\ldots,V_{k}$. Therefore, we have a polynomial (in $n$) number of possible solutions for Maximum Total Transitivity Decision Problem when the total transitivity of the graph is constant. Also, note that we can verify whether a vertex partition is a total transitive partition or not in polynomial time. Therefore, when the total transitivity of the graph is constant, we can solve Maximum Total Transitivity Decision Problem in polynomial time. Hence, we have the following corollary.
Corollary 10.
The Maximum Total Transitivity Decision Problem is a fixed parameter tractable when parameterized by the size of the total transitive partition.
In the Remark 1, we have already shown that the difference between transitivity and total transitivity can be arbitrarily large. Now we show another non-trivial example where the difference between these two parameters is arbitrarily large. Let $G=(S\cup K,E)$ be a split graph, where $S$ and $K$ are an independent set and clique of $G$, respectively. Also, assume that $K$ is the maximum clique of $G$, that is, $\omega(G)=|K|$. Furthermore, $|K|=|S|$, and each vertex of $S$ is adjacent to exactly one vertex of $K$ (see Figure 1).
Figure 1: Example of a split graph where $Tr_{t}(G)=1$ and $Tr(G)=\omega(G)+1=6$
First, we show that for the above split graph, $Tr_{t}(G)=1$. Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of the above split graph $G$. Since $deg(s)=1$ for all $s\in S$, $s$ cannot be in $V_{p}$ for $p\geq 2$. Moreover, $\pi$ is a total transitive, which implies that to dominate $s\in S$, there must be a vertex from $K$ in $V_{1}$. From the above example graph, it is clear that to dominate vertices from $S$, all the vertices of $K$ must be in $V_{1}$. Hence, $Tr_{t}(G)=1$. Again, we know from [18], $Tr(G)=\omega(G)+1$, as $S$ dominates $K$. Therefore, for large $\omega(G)$, the difference $Tr(G)-Tr_{t}(G)=\omega(G)+1-1=\omega(G)$ is arbitrarily large.
Next, we prove the existence of graphs with transitivity is one more than total transitivity. We define a special tree, namely total complete minimum broadcast tree of order $k$ ($tcmbt_{k}$). Apart from this property, the $tcmbt_{k}$ possesses some important properties. The recursive definition of $tcmbt_{k}$ is as follows:
Definition 11.
A total complete minimum broadcast tree of order $k$ $(tcmbt_{k})$ is defined recursively as follows:
  1. 1.
    The tcmbt of order $1$ $(tcmbt_{k})$ is a rooted tree, whose root has degree $1$ and adjacent to a vertex of degree $1$.
  2. 2.
    The tcmbt of order $k$ $(tcmbt_{k})$ is a rooted tree, whose root has degree $k$ and is adjacent to vertices $\{v_{1},v_{2},\ldots,v_{k}\}$, such that $v_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1. Moreover, the vertex $v_{k}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1.
Figure 2: $tcmbt$ of order upto $3$ along with its unique total transitive partition
Figure 2, illustrates the $tcmbt$ of order upto $3$. Note that from the definition, it follows that the root of $tcmbt_{k}$ and exactly one of its children are the only vertices that can be in $V_{k}$, in a total transitive partition of order $k$. Because of the acyclic property, none of the vertices in $V_{i}$ can dominate more than one vertex from $V_{j}$ with $j>i$. Therefore, the $tcmbt_{k}$ is the minimum tree (in terms of vertices) with total transitivity as $k$, and it has a unique total transitive partition of size $k$ (refer Figure 2). This property of $tcmbt_{k}$ helps us to use these graphs as a gadget in the NP-completeness proofs in Section 4. We show that if a graph $G$ is a $tcmbt_{k}$, then $Tr_{t}(G)=k$ and $Tr(G)=k+1$. For that, we have the following lemma.
Lemma 12.
Let $T$ be a $tcmbt_{k}$. Then $Tr_{t}(T)=k$ and $Tr(T)=k+1$.
Proof.
First, we show that $Tr_{t}(T)=k$. For this, we use induction on the order of $tcmbt$. For the base case, that is, when $k=1$, by the definition of $tcmbt$, $T$ is a $P_{2}$. Therefore, $Tr_{t}(P_{2})=1$. Let $T$ be a $tcmbt_{k}$. By the definition, we have $T$ is a rooted tree, whose root has degree $k$ and is adjacent to vertices $\{v_{1},v_{2},\ldots,v_{k}\}$, such that $v_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1. Moreover, the vertex $v_{k}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1. Let $c$ be the root of $T$ and $c^{\prime}$ be its child such that $c^{\prime}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1. By the induction hypothesis, $Tr_{t}(T_{v_{j}})=Tr_{t}(T_{u_{j}})=j$, for all $11\leq i\leq k-$ 1, where $T_{v}$ is a subtree of $T-\{c\}$, rooted at $v$ (see the $tcmbt_{3}$ in Figure 2). Now $T-\{c,c^{\prime}\}$ is a disjoint union of trees $T_{v_{j}},T_{u_{j}}$, for all $11\leq j\leq k-$ 1, and $k-1=\max\{Tr_{t}(T_{v_{j}})|1\leq j\leq k-1\}=\max\{Tr_{t}(T_{u_{j}})|1\leq j\leq k-1\}$. As total transitivity depends only on the maximum total transitivity of a component of a disconnected graph (PropositionLABEL:disjoint_union_total_transitivity). Therefore, $Tr_{t}(T-\{c,c^{\prime}\})$ is $k-1$. Moreover, we can construct a total transitive partition, say $\pi=\{V_{1},V_{2},\ldots,V_{k-1}\}$, of $T-\{c,c^{\prime}\}$ recursively such that $\{v_{i},u_{i}\}\in V_{i}$ for all $11\leq i\leq k-$ 1. Now clearly, $\pi^{\prime}=\pi\cup\{c,c^{\prime}\}$ forms a total transitive partition of $T$. Therefore, $Tr_{t}(T)\geq k$. Moreover, by the Proposition 1, we know that $Tr_{(}T)\leq\Delta(G)=k$. Hence, $Tr_{t}(T)=k$. The algorithm described in [14] shows that $Tr(T)=k+1$. Therefore, $Tr_{t}(T)=k$ and $Tr(T)=k+1$ for a $tcmbt_{k}$. ∎
3 Total Transitivity in split graphs
In this section, we show that total transitivity of a split graph can be between $1$ and $\omega(G)-1$, where $\omega(G)$ is the size of a maximum clique in $G$. Then, we characterize split graphs with total transitivity of $1$ and $\omega(G)-1$. Later, we show some necessary conditions for $Tr_{t}(G)=p$, $11\leq p\leq\omega(G)-$ 1. First, we prove that total transitivity of a split graph $G$ cannot be greater than $\omega(G)-1$.
Lemma 13.
Let $G=(S\cup K,E)$ be a split graph, where $S$ and $K$ are an independent set and clique of $G$, respectively. Also, assume that $K$ is the maximum clique of $G$, that is, $\omega(G)=|K|$. Then $11\leq Tr_{t}(G)\leq\omega(G)-$ 1.
Proof.
Clearly, $Tr_{t}(G)\geq 1$. Also, from the definition of total transitivity, we know that $Tr_{t}(G)\leq Tr(G)$. Again, from [18], we know $Tr(G)\leq\omega(G)+1$. Therefore, for a split graph, $Tr_{t}(G)\leq\omega(G)+1$. Now we show that $Tr_{t}(G)$ cannot be $\omega(G)+1$ or $\omega(G)$. Let $Tr_{t}(G)=\omega(G)+1$ and $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)+1}\}$ be a total transitive partition of $G$. Since $|K|=\omega(G)$, there exists at least one set in $\pi$, say $V_{i}$, such that $V_{i}$ contains only vertices from $S$. Note that, in this case, $V_{i}$ cannot dominate $V_{i}$ as $S$ is an independent set of $G$. Therefore, we have a contradiction. Hence, $Tr_{t}(G)<\omega(G)+1$.
[…]
Next, we characterize split graphs with $Tr_{t}(G)=1$. A vertex is said to be a leaf vertex if its degree is $1$, and the vertex that is adjacent to a leaf vertex is said to be the support vertex of the leaf vertex.
Theorem 14.
Let $G=(S\cup K,E)$ be a split graph, where $S$ and $K$ are the independent set and clique of $G$, respectively. Also, assume that $K$ is the maximum clique of $G$, that is, $\omega(G)=|K|$. Then $Tr_{t}(G)=1$ if and only if $G$ satisfies one of the following conditions.
  1. (a)
    The $deg(s)=1$ for all $s\in S$ and $K$ has at most vertex without a neighbour in $S$.
  2. (b)
    For all $s\in S$ with $deg(s)\geq 2$, each vertex from $N(s)$ is a support vertex, and $K$ has at most one vertex without a neighbour in $S$.
Proof.
Let $G$ satisfy the condition $(a)$. We show that $Tr_{t}(G)=1$. Assume $\pi=\{V_{1},V_{2},\ldots,V_{t}\}$ is a total transitive partition of $G$. Since $deg(s)=1$ for all $s\in S$, no vertices from $S$ can be in $V_{p}$ for $p\geq 2$. To dominate vertices from $S$, $V_{1}$ contains all vertices from $K$ except at most one. So $t$ cannot be more than $2$, and $V_{2}$ can contain at most one vertex from $K$, which is a contradiction as $\pi$ is a total transitive partition. Hence, $Tr_{t}(G)=1$.
For the condition $(b)$, let $\pi=\{V_{1},V_{2},\ldots,V_{t}\}$ is a total transitive partition of $G$. Clearly, if the $deg(s)=1$ for some $s\in S$, it must belong to $V_{1}$, so assume there exists $s^{\prime}\in S$ and $deg(s^{\prime})\geq 2$. By the condition $(b)$, each vertex from $N(s^{\prime})$ is a support vertex. Therefore, each vertex from $N(s^{\prime})$ must be in $V_{1}$, and hence $s^{\prime}$ cannot be in $V_{p}$ for $p\geq 2$. From above, we have all vertices must be in $V_{1}$, and as $K$ has at most one vertex without a neighbour in $S$, $V_{p}$ must be empty from $p\geq 2$. Hence, $Tr_{t}(G)=1$.
Conversely, assume that $Tr_{t}(G)=1$. We divide our proof into two cases:
Case 1.
The $deg(s)=1$ for all $s\in S$
We show that $K$ has at most one vertex without a neighbour in $S$. If $K$ has more than one vertex without a neighbour in $S$, then we show a contradiction. Let $\{k_{1},k_{2}\}\subseteq K$ such that $k_{1},k_{2}$ has no neighbour in $S$. Let us consider a vertex partition $\{V_{1},V_{2}\}$ of $G$ as follows: $V_{2}=\{k_{1},k_{2}\}$, and $V_{1}$ contains all the other vertices of $G$. Clearly, $\{V_{1},V_{2}\}$ is a total transitive partition of $G$ with size $2$, which implies that $Tr_{t}(G)\geq 2$, a contradiction as $Tr_{t}(G)=1$. Hence, $K$ has at most one vertex without a neighbour in $S$, and $G$ satisfies condition $(a)$.
Case 2.
There exists $s\in S$ with $deg(s)\geq 2$
Let $s\in S$ and $deg(s)\geq 2$. Also, let $N(s)=\{k_{1},k_{2},\ldots,k_{q}\}$, where $k_{i}\in K$. If there exists a $j$ such that $k_{j}$ is not a support vertex, then $\{V_{1},V_{2}\}$, where $V_{2}=\{s,k_{j}\}$ and $V_{1}$ contains all the other vertices of $G$, form a total transitive partition of $G$ of size $2$. Therefore, we have a contradiction as $Tr_{t}(G)=1$. So, all the vertices from $N(s)=\{k_{1},k_{2},\ldots,k_{q}\}$ are support vertex. Again, if $K$ has more than one vertex without a neighbour in $S$, similar to Case 1, we can show a contradiction. Hence, in this case, for all $s\in S$ with $deg(s)\geq 2$, each vertex from $N(s)$ is a support vertex, and $K$ has at most one vertex without a neighbour in $S$; that is, $G$ satisfies condition $(b)$. ∎
Next, we characterize split graphs with $Tr_{t}(G)=\omega(G)-1$. Before characterizing the split graphs with $Tr_{t}(G)=\omega(G)-1$, we have the following observations for any $p$, $11\leq p\leq q-$ 1, where $q=\omega(G)$.
Observation 15.
Let $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of a split graph $G$. Then, each $V_{i}$ must contain at least one vertex from $K$.
Observation 16.
Let $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of a split graph $G$. If $V_{j}$, $j\neq 1$ contains more than one vertices from $S$, then there exists $\pi^{\prime}=\{V_{1}^{\prime},V_{2}^{\prime},\ldots,V_{p}^{\prime}\}$ such that $|V_{j}^{\prime}\cap S|\leq 1$ for all $j\neq 1$. Moreover, when $p=\omega(G)-1$, there exists a total transitive partition of $G$, say $\pi^{\prime\prime}=\{V_{1}^{\prime\prime},V_{2}^{\prime\prime},\ldots,V_{p}^{\prime\prime}\}$ such that $|V_{i}|=2$ for all $2\leq i\leq p$.
Observation 17.
Let $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of $G$ and $x\in S$ such that $deg(x)=1$. Then $x$ and its support vertex, say $y\in K$, cannot be in $V_{t}$ for $t\geq 2$.
Observation 18.
Let $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of $G$. Then there must exist a subset of $K$, say $K^{\prime}$, such that $K^{\prime}$ dominates $S$, and all the vertices of $K^{\prime}$ must be in $V_{1}$. Moreover, $11\leq|K^{\prime}|\leq q-p+$ 1, that is, the size of the $K^{\prime}$ is at most $q-p+1$.
From Observation 18, we have if $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ is a total transitive partition of $G$. Then there must exist a subset of $K$, say $K^{\prime}$, such that $K^{\prime}$ dominates $S$, and all the vertices of $K^{\prime}$ must be in $V_{1}$. Moreover, $11\leq|K^{\prime}|\leq q-p+$ 1, that is, the size of the $K^{\prime}$ is at most $q-p+1$. Let us define the minimum cardinality of $K^{\prime}$ as S domination number by K and denote it by $dom_{K}(S)$ and $K^{\prime}$ with minimum cardinality as $dom_{K}(S)$-set.
Let $G$ be a split graph, and $Tr_{t}(G)=p$ for some $11\leq p\leq\omega(G)-$ 1. Also, assume that $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of $G$ with size $p$. Then we say $dom_{K}(S)$-set exists if there exists a subset $K^{\prime}$ of $K$, such that $K^{\prime}$ dominates $S$, vertices of $K^{\prime}$ are in $V_{1}$ and $11\leq|K^{\prime}|\leq q-p+$ 1, where $q=\omega(G)$.
In the following theorem, we characterize the split graphs with $Tr_{t}(G)=\omega(G)-1$.
Theorem 19.
Let $G=(S\cup K,E)$ be a split graph, where $S$ and $K$ are the independent set and clique of $G$, respectively. Also, assume that $K$ is the maximum clique of $G$, that is, $\omega(G)=|K|$. Then $Tr_{t}(G)=\omega(G)-1$ if and only if $G$ satisfies one of the following conditions.
  1. (a)
    $dom_{K}(S)$-set exists and $dom_{K}(S)=1$. Then there exist $\{s_{1},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $1\leq j\leq\alpha$, where $\alpha=\omega(G)-2$ Also, for some pair of vertices $x,y$, we have an ordering of the vertices of $K\setminus\{k_{1},x,y\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1}\}$ is a $dom_{K}(S)$-set.
  2. (b)
    $dom_{K}(S)$-set exists and $dom_{K}(S)=2$. Then there exist $\{s_{2},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $2\leq j\leq\alpha$, where $\alpha=\omega(G)-1$ Also, we have an ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1},k_{1}^{\prime}\}$ is a $dom_{K}(S)$-set.
Proof.
Let $G$ be a split graph, and $G$ satisfies the condition $(a)$. We show that $Tr_{t}(G)=\omega(G)-1$. From the condition $(a)$, there exist $\{s_{1},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $1\leq j\leq\alpha$, where $\alpha=\omega(G)-2$ Also, for some pair of vertices $x,y$, we have an ordering of the vertices of $K\setminus\{k_{1},x,y\}$, say $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1}\}$ is a $dom_{K}(S)$-set. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)-1}\}$ of $G$ as follows: $\{k_{1},s_{1}\}\subseteq V_{1}$, $V_{j}=\{s_{j},k_{j}\}$, for all $2\leq j\leq\alpha$, $V_{\omega(G)-1}=\{x,y\}$, and all the other vertices of $G$ in $V_{1}$. We show that $\pi$ is a total transitive partition. Since $\{k_{1}\}$ is a $dom_{K}(S)$-set and $K$ is a clique, $V_{1}$ dominates $V_{i}$ for all $i\geq 1$. Again, from condition $\{k_{2},k_{3},\ldots,k_{j}\}\subseteq N(s_{j})$, we have $V_{s}$ dominates $V_{t}$ for all $22\leq s\leq t\leq\omega(G)-$ 2. Furthermore, as $V_{\omega(G)-1}=\{x,y\}$ and $K$ is a clique, $V_{s}$ dominates $V_{\omega(G)-1}$ for all $11\leq s\leq\omega(G)-$ 1. Therefore, $\pi$ is a total transitive partition of size $\omega(G)-1$, which implies $Tr_{t}(G)\geq\omega(G)-1$. Moreover, from Lemma13, we have $Tr_{t}(G)\leq\omega(G)-1$. Hence, if $G$ satisfies condition $(a)$, $Tr_{t}(G)=\omega(G)-1$.
Now assume $G$ be a split graph and $G$ satisfies the condition $(b)$. We show that $Tr_{t}(G)=\omega(G)-1$. From the condition $(b)$, there exist $\{s_{2},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $2\leq j\leq\alpha$, where $\alpha=\omega(G)-1$ Also, we have an ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1},k_{1}^{\prime}\}$ is a $dom_{K}(S)$-set. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)-1}\}$ of $G$ as follows: $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$, $V_{j}=\{s_{j},k_{j}\}$, for all $2\leq j\leq\alpha$, and all the other vertices of $G$ in $V_{1}$. We show that $\pi$ is a total transitive partition. Since $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$ is a $dom_{K}(S)$-set and $K$ is a clique, $V_{1}$ dominates $V_{i}$ for all $i\geq 1$. Again, from condition $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, we have $V_{s}$ dominates $V_{t}$ for all $12\leq s\leq t\leq\omega(G)-$ 1. Therefore, $\pi$ is a total transitive partition of size $\omega(G)-1$, which implies $Tr_{t}(G)\geq\omega(G)-1$. Moreover, from Lemma13, we have $Tr_{t}(G)\leq\omega(G)-1$. Hence, if $G$ satisfies condition $(b)$, $Tr_{t}(G)=\omega(G)-1$.
Therefore, if $G$ is a split graph and satisfies either $(a)$ or $(b)$ then $Tr_{t}(G)=\omega(G)-1$.
Now we prove the converse part of the theorem. Let $Tr_{t}(G)=\omega(G)-1$. Then $dom_{K}(S)$-set exists and $21\leq|dom_{K}(S)|\leq\omega(G)-\{\omega(G)-1\}+1=$ 2. Based on $|dom_{K}(S)|=1$ or $|dom_{K}(S)|=2$, we divide our proof into two following cases:
Case 1.
$|dom_{K}(S)|=1$
By the Observation17, we can take a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ of $G$ such that $|V_{i}|=2$ for all $2\leq i\leq p$, where $p=\omega(G)-1$. Let $\{k_{1}\}\subseteq K$ be a $dom_{K}(S)$-set. Again from Observation15, we know that each $V_{i}$ must contain a vertex from $K$. Therefore, exactly one set, say $V_{r}$ from $\pi$, contains two vertices from $K$. First, assume $V_{r}$ is not $V_{1}$, then $V_{1}$ contains exactly one vertex from $K$ and at least one vertex from $S$. Also, each set $V_{2},\ldots,V_{r-1},V_{r+1},\ldots,V_{p}$ contains exactly one vertex from $K$ and exactly one vertex from $S$. Without loss of generality, assume $\{s_{1},k_{1}\}\subseteq V_{1}$ and $V_{j}=\{s_{j},k_{j}\}$ for all all $1\leq j\leq r-1,r+1\leq j\leq p$ and $V_{r}=\{k_{r},k_{r}^{\prime}\}$. Since $\pi$ is a total transitive partition of $G$ the $deg(s_{j})\geq j$ for all $1\leq j\leq p$ and $j\neq r$. Moreover, $\{k_{1},k_{2},\ldots,k_{j}\}\subseteq N(s_{j})$ for all $11\leq j\leq r-$ 1 and $\{k_{1},k_{2},\ldots,k_{r-1},k_{r+1},\ldots,k_{j}\}\subseteq N(s_{j})$ for all $r+1\leq j\leq p$. Clearly, we have the required ordering of the vertices of $K\setminus\{k_{1},k_{r},k_{r}^{\prime}\}$. Similarly, we can prove that $G$ satisfies the condition $(a)$ when $V_{r}$ is $V_{1}$.
Case 2.
$|dom_{K}(S)|=2$
In this case, we can also assume there exists a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ of $G$ such that $|V_{i}|=2$ for all $2\leq i\leq p$, where $p=\omega(G)-1$. Let $\{k_{1},k_{1}^{\prime}\}\subseteq K$ be a $dom_{K}(S)$-set. From Observation15 we know that each $V_{i}$ must contains a vertex from $K$ and by definition of $dom_{K}(S)$-set $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$. So, each $V_{i}$ contains exactly one vertex from $K$ and exactly one vertex from $S$, for all $2\leq i\leq p$. Without loss of generality assume $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$ and $V_{i}=\{s_{i},k_{i}\}$ for all all $1\leq i\leq p$. Since $\pi$ is a total transitive partition of $G$ the $deg(s_{i})\geq j$ for all $2\leq i\leq p$. Moreover, $\{k_{2},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $1\leq i\leq p$. Clearly, we have the required ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$.
Hence, $Tr_{t}(G)=\omega(G)-1$ if and only if $G$ satisfies one of the above conditions. ∎
From the Lemma 13, we have $11\leq Tr_{t}(G)\leq\omega(G)-$ 1. In the above, we already discussed the characterization of split graphs when $Tr_{t}(G)=1$, and $Tr_{t}(G)=\omega(G)-1$. Now, we show some necessary conditions for $Tr_{t}(G)=p$, $11\leq p\leq q-$ 1, where $q=\omega(G)$.
Theorem 20.
Let $G=(S\cup K,E)$ be a split graph, where $S$ and $K$ are the independent set and clique of $G$, respectively. Also, assume that $K$ is the maximum clique of $G$, that is, $\omega(G)=|K|$. If $Tr_{t}(G)=p$, then $G$ satisfies one of the following conditions.
  1. (a)
    $dom_{K}(S)$-set exists and $dom_{K}(S)=1$
    1. (i)
      $|S|\geq 2p-q$
    2. (ii)
      There exists $\{s_{1},s_{2},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $1\leq j\leq\alpha$, where $\alpha=2p-q$
  2. (b)
    $dom_{K}(S)$-set exists and $dom_{K}(S)\geq 2$
    1. (i)
      $|S|\geq 2p-q+l_{min}-2$, where $l_{min}=|dom_{K}(S)|$
    2. (ii)
      There exists $\{s_{2},\ldots,s_{\beta}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $2\leq j\leq\beta$, where $\beta=2p-q+l_{min}-2$
Proof.
Let $Tr_{t}(G)=p$ and $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of $G$ with size $p$. By the Observation18, $dom_{K}(S)$-set exists and $11\leq|dom_{K}(S)|\leq p-q+$ 1. We divide our proof into two cases:
Case 1.
$|dom_{K}(S)|=1$
As $\pi$ is a total transitive partition, size each $|V_{i}|\geq 2$. Also, by Observation15, each $V_{i}$ must contain a vertex from $K$. Now it is given that $Tr_{t}(G)=p$ and $|K|=q$, so each $p-(q-p)=2p-q$ set from $\pi$ must contain at least one vertex from $S$. Therefore, $|S|\geq p-(q-p)=2p-q$ $(a(i))$. Again, by the Observation16, each $|V_{i}\cap S|\leq 2$. Let $\{s_{i_{1}},s_{i_{2}},\ldots,s_{i_{\alpha}}\}\subseteq S$, where $\alpha=2p-q$ such that $s_{i_{j}}\in V_{i_{j}}$. Since $\pi$ is a total transitive partition, $deg(s_{i_{j}})\geq i_{j}\geq j$ $(a(ii))$.
Case 2.
$|dom_{K}(S)|=2$
In this case, let $l_{min}=|dom_{K}(S)|$. By the definition of $dom_{K}(S)$, the vertices are in $V_{1}$. As we know from the Observation15, each $V_{i}$ must contain a vertex from $K$, so each $(p-1)-[q-\{(p-1)+l_{min}\}]=2p-q+l_{min}-2$ set from $\pi$ must contain at least one vertex from $S$. Therefore, $|S|\geq 2p-q+l_{min}-2$ $(b(i))$. Again, by Observation16, each $|V_{i}\cap S|\leq 2$. Let $\{s_{i_{1}},s_{i_{2}},\ldots,s_{i_{\beta}}\}\subseteq S$, where $\beta=2p-q+l_{min}-2$ such that $s_{i_{j}}\in V_{i_{j}}$. Also, note that $i_{j}\geq 2$. Since $\pi$ is a total transitive partition, $deg(s_{i_{j}})\geq i_{j}\geq j$ $(b(ii))$.
Hence, if $Tr_{t}(G)=p$, $G$ satisfies one of the above conditions. ∎
4 NP-completeness
In this section, we show that the Maximum Total Transitivity Decision Problem is NP-complete for a bipartite graph. Clearly, this problem is in NP. We prove the NP-hardness of this problem by showing a polynomial-time reduction from Proper $3$-Coloring Decision Problem in graphs to the MtotTDP. A proper $3$-colring of a graph $G=(V,E)$ is a function $g:V\rightarrow\{1,2,3\}$, such that for every edge $uv\in E$, $g(u)\not=g(v)$. The proper $3$-Coloring Decision Problem is defined as follows:
Proper $3$-Coloring Decision Problem (P3333CDP) Instance: A graph $G=(V,E)$ Question: Does there exist a proper $3$-coloring of $V$?
The P3333CDP is known to be NP-complete for graphs [7]. Given an instance $G=(V,E)$ of P3333CDP, we construct an instance of Maximum Total Transitivity Decision Problem, that is, a graph $G^{\prime}=(V^{\prime},E^{\prime})$ and a positive integer $k$.
Figure 3: Construction of $G^{\prime}$ from $G$ where $m=3$ and $T$ represents a $tcmbt_{3}$
The construction is as follows. Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$ and $E=\{e_{1},e_{2},\ldots,e_{m}\}$. For each vertex $v_{i}\in V$, we consider two $tcmbt_{3}$ with $v_{i}$ and $v_{i}^{\prime}$ as their roots, respectively. Similarly, for each edge $e_{i}\in E$, we consider another two $tcmbt_{3}$ with $v_{e_{i}}$ and $v_{e_{i}}^{\prime}$ as their roots, respectively. Further, corresponding to every edge $e_{j}\in E$, we take two vertices $e_{j},e_{j}^{\prime}$ in $G^{\prime}$. Also, we take another two vertices $e$ and $e^{\prime}$ in $G^{\prime}$. Let $A=\{e_{1},e_{2},\ldots,e_{m},e\}$ and $B=\{e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime}\}$. We construct a complete bipartite graph with vertex set $A\cup B$. Now, for each edge $e_{t}=v_{i}v_{j}\in E$, we join the edges $v_{i}e_{t}$, $v_{j}e_{t}$, $v_{e_{t}}e_{t}$, $v_{i}^{\prime}e_{t}^{\prime}$, $v_{j}^{\prime}e_{t}^{\prime}$, and $v_{e_{t}}^{\prime}e_{t}^{\prime}$ in $G^{\prime}$. Next, we consider six $tcmbt_{3}$ with roots $v_{a},v_{a}^{\prime},v_{e},v_{e}^{\prime},v_{b},v_{b}^{\prime}$. Finally, we make $\{v_{a},v_{e},v_{b}\}$, $\{v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$ adjacent to the vertex $e$, and $e^{\prime}$, respectively, and we set $k=m+4$. In this construction, we can verify that the graph $G^{\prime}=(V^{\prime},E^{\prime})$ consists of $11640m+38n+$ 116 vertices and $m^{2}+42m+34n+104$ edges. The construction is illustrated in Figure3.
Next, we show that $G$ has a proper $3$-coloring if and only if $G^{\prime}$ has a total transitive partition of size $k$. First, we show the forward direction of this statement in the following lemma.
Lemma 21.
If $G=(V,E)$ has a proper $3$-coloring, then $G^{\prime}=(V^{\prime},E^{\prime})$ has a total transitive partition of size $k$.
Proof.
Let $g$ be a proper $3$-coloring of $G$. Based on $g$, let us consider a vertex partition, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, of $G^{\prime}$ as follows: for each $v_{i}\in V$, we put the corresponding vertex $v_{i},v_{i}^{\prime}$ in $V_{q}$, if $g(v_{i})=q$. For each edge $e_{t}=v_{i}v_{j}\in E$, we put the vertex $v_{e_{t}},v_{e_{t}}^{\prime}\in V_{q}$, if $g(v_{i})\neq q$ and $g(v_{j})\neq q$. Further, we put $v_{a},v_{a}^{\prime}\in V_{3}$, $v_{e},v_{e}^{\prime}\in V_{2}$, and $v_{b},v_{b}^{\prime}\in V_{1}$. We put the other vertices of $tcmbt_{3}$ according to Figure 4, based on the position of the root vertex in the partition. Finally, for $1\leq j\leq m$, we put $e_{j},e_{j}^{\prime}\in V_{3+j}$ and $e,e^{\prime}\in V_{m+4}$.
Figure 4: Partition of $tcmbt_{3}$
Since the vertices of $A\cup B$ induces a complete bipartite graph, $V_{i}$ dominates $V_{j}$ for $4\leq i\leq j\leq k$. Also, for each $i=1,2,3$, every vertex of $A\cup B$ is adjacent to two vertices of $V_{i}$. Therefore, for each $i=1,2,3$, $V_{i}$ dominates $V_{t}$, for all $t>3$. Finally, from Figure4, it is evident that $V_{i}$ dominates $V_{j}$ for $1\leq i\leq j\leq 3$. Hence, $\pi$ is a total transitive partition of $G^{\prime}$ of size $k$. Therefore, if $G$ has a proper $3$-coloring, then $G^{\prime}$ has a total transitive partition of size $k$. ∎
Next, we show the converse of the statement is true, that is, if $G^{\prime}$ has a total transitive partition of size $k$ then $G$ has a proper $3$-coloring. For this, first, we argue the existence of a total transitive partition of $G^{\prime}$, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that the sets $V_{1},V_{2},$ and $V_{3}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$. If such a total transitive partition exists, then we can define a $3$-coloring function $g$ from $V(G)$ to $\{1,2,3\}$, using the partition $\pi$ as follows: $g(v)=l$, if and only if $v\in V_{l}$. In the following claim, we show the existence of such a total transitive partition of $G^{\prime}$.
Claim 22.
Let  $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of $G^{\prime}$ of size $k$ such that $|V_{k}|=2$. Then the sets $V_{4},V_{5},\ldots V_{k}$ contain only vertices from $A\cup B$, and the sets $V_{1},V_{2}$, and $V_{3}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.
Proof.
First note that the vertices $\{v_{a},v_{e},v_{b}\}$, which are adjacent to $e$, must belong to $V_{1},V_{2}$ or $V_{3}$. As the degree of each vertex from $\{v_{a},v_{e},v_{b}\}$ is $4$, they can be at most in $V_{4}$, but in this case, $e$ can be either in $V_{3}$ or $V_{4}$. Furthermore, if $e$ is in $V_{3}$, there is no neighbour in $V_{4}$ to dominate vertices from $\{v_{a},v_{e},v_{b}\}$ and similarly if $e$ is in $V_{4}$, then here is no neighbour in $V_{3}$ to dominate vertices from $\{v_{a},v_{e},v_{b}\}$. Similarly, the vertices of $\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$, which are adjacent to $e^{\prime}$ cannot be in $V_{p}$ for any $p\geq 4$. Further, since the degree of every vertex other than the vertices of $A\cup B$ is less than $m+4$, $V_{k}$ can contain two vertices only from $A\cup B=\{e_{1},e_{2},\ldots,e_{m},e,e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime}\}$. In this proof, we show that to dominate the vertices of $A\cup B$ which belong to $V_{k}$, the other vertices of $A\cup B$ must be in $V_{4},V_{5},\ldots,V_{k-1}$. This, in turn, would imply that the sets $V_{1},V_{2}$, and $V_{3}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$. The proof is divided into four cases based on whether $e\in V_{k}$ or not and $e^{\prime}\in V_{k}$ or not.
Case 1.
$e\in V_{k}$ and $e^{\prime}\in V_{k}$
As discussed earlier, the vertices $\{v_{a},v_{e},v_{b}\}$ and $\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$. Since $e\in V_{k}$ to dominate $e$, each set in $\{V_{1},\ldots,V_{k-1},V_{k}\}$ must contain at least one vertices from $N_{G^{\prime}}(e)=\{e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime},v_{a},v_{e},v_{b}\}$. Since $e$ is adjacent to exactly $k=m+4$ vertices, each $V_{i},1\leq i\leq k$ contains exactly one vertex from $N_{G^{\prime}}(e)$. Since $\{v_{a},v_{e},v_{b}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$ it follows that vertices from $N_{G^{\prime}}(e)\setminus\{v_{a},v_{e},v_{b}\}$ belong to $V_{4},V_{5},\ldots V_{k}$. Similarly, as $e^{\prime}\in V_{k}$, to dominate $e^{\prime}$, each set in $\{V_{1},V_{2},\ldots,V_{k-1},V_{k}\}$ must contain at least one vertex from $N_{G^{\prime}}(e^{\prime})=\{e_{1},e_{2},\ldots,e,v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$. Since $e^{\prime}$ is adjacent to exactly $k=m+4$ vertices, each $V_{i},1\leq i\leq k$ contains exactly one vertex from $N_{G^{\prime}}(e^{\prime})$. Similarly, as before, since $\{v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$, it follows that exactly three vertices from $N_{G^{\prime}}(e^{\prime})\setminus\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$ belong to $V_{4},V_{5},\ldots V_{k}$. Hence, the vertices of $A\cup B$ belong to $\{V_{4},V_{5},\ldots V_{k}\}$.
Clearly, none of the vertices from the set $\{v_{1},\ldots,v_{n},v^{\prime}_{1},\ldots,v^{\prime}_{n}\}$ and the set $\{v_{e_{1}},\ldots,v_{e_{m}},v^{\prime}_{e_{1}},\ldots,v^{\prime}_{e_{m}}\}$ can belong to $V_{p}$ for $p\geq 4$ because otherwise there exists a vertex in $A\cup B$ which belongs to $V_{3}$. Since the degree of every other vertices is at most $3$, they cannot belong to $V_{p}$ for $p\geq 4$. Hence, $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$ and $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.
Case 2.
$e\in V_{k}$ and $e^{\prime}\notin V_{k}$
Using similar arguments as in the previous case, we know that vertices from $\{v_{a},v_{b},v_{e}\}$ belong to $V_{p}$ for $1\leq p\leq 3$ and the vertices from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1},V_{k}\}$ to dominate $e$. Moreover, every $V_{i}$ for $4\leq i\leq k$ contains exactly one vertex from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$. Since the vertices from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1},V_{k}\}$, no vertex from $\{v^{\prime}_{1},\ldots,v^{\prime}_{n},v^{\prime}_{e_{1}},\ldots,v^{\prime}_{e_{m}},v^{\prime}_{a},v^{\prime}_{b},v^{\prime}_{e}\}$ can be in $V_{p}$ for $p\geq 4$. As $e^{\prime}\notin V_{k}$, there exist a vertex from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m}\}$, say $e^{\prime}_{r}$ such that $e^{\prime}_{r}\in V_{k}$. To dominate $e^{\prime}_{r}$, the vertices from $\{e_{1},e_{2},\ldots,e_{m},e\}$ belong to $\{V_{4},V_{5},\ldots,V_{k}\}$. With similar arguments as in the previous case, we know that every vertex of $\{v_{1},v_{2},\ldots,v_{n},v_{e_{1}},v_{e_{2}},\ldots,v_{e_{m}}\}$ belong to $V_{p}$ for $1\leq p\leq 3$. Hence, $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$ and $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.
Case 3.
$e\notin V_{k}$ and $e^{\prime}\in V_{k}$
Using similar arguments as in Case 2, we can show that $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$ and $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.
Case 4.
$e\notin V_{k}$ and $e^{\prime}\notin V_{k}$
Since the degree of each vertex of $V^{\prime}\setminus(A\cup B)$ is at most $k-1=m+3$, they cannot belong to $V_{k}$, that is, only vertices from $A\cup B$ can be in $V_{k}$. Without loss of generality, we can assume that $e_{1}\in V_{k}$ and $e^{\prime}_{s}\in V_{k}$ in the total transitive partition of $G^{\prime}$ of size $k$. Also, let $e_{1}$ be the edge between $v_{1}$ and $v_{2}$ in $G$ and in the total transitive partition of $G^{\prime}$ of size $k$, $v_{1}\in V_{l}$ and $v_{2}\in V_{t}$ where $l\geq t$. Next, we show that $l\leq 3$. As $v_{1}\in V_{l}$ and $\pi$ is a total transitive partition of $G$ implies that each of $V_{3},V_{4},\ldots,V_{l}$ contains at least one vertex from $A\setminus\{e_{1},e\}$. On the other hand, as $e_{s}^{\prime}\in V_{k}$, each of $\{V_{l+1},\ldots,V_{k-1}\}$ also contains at least one vertex from $A\setminus\{e_{1}\}$. This implies that each of the $(m+1)$ partitions, $\{V_{3},V_{4},\ldots,V_{k-1}\}$, contains one vertex from the set of $m$ vertices, $A\setminus\{e_{1}\}$, which is impossible. Hence, $l\leq 3$. With a similar argument, we can show that $v_{e_{1}}$ cannot be in $V_{p}$ for $p\geq 4$. Therefore, each $\{V_{3},V_{4},\ldots,V_{k}\}$ contains exactly one vertex from $B$. Since the vertices of $B$ belong to $\{V_{4},V_{5},\ldots,V_{k}\}$, the vertices of $\{v^{\prime}_{1},v^{\prime}_{2},\ldots,v^{\prime}_{n},v^{\prime}_{e_{1}},v^{\prime}_{e_{2}},\ldots,v^{\prime}_{e_{m}},v^{\prime}_{a},v^{\prime}_{b},v^{\prime}_{e}\}$ belong to some partition $V_{p}$ for some $p\leq 3$. As $e^{\prime}_{s}\in V_{k}$ and $\pi$ is a total transitive partition of $G$ to dominate $e^{\prime}_{s}$, the vertices from $A\setminus\{e_{1}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1}\}$. With similar arguments as in Case 1, we know that each vertex of $\{v_{1},v_{2},\ldots,v_{n},v_{e_{1}},v_{e_{2}},\ldots,v_{e_{m}},v_{a},v_{b},v_{e}\}$ belongs to $V_{p}$ for some $p\leq 3$. Hence, $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$ and $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.
Therefore, for all the cases, the partitions $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$ and the partitions $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$. ∎
Now we prove $G$ uses $3$-colors, which is a proper coloring in the following lemma.
Lemma 23.
If $G^{\prime}$ has a total transitive partition of size $k$, then $G$ has a proper $3$-coloring.
Proof.
Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of $G^{\prime}$ of size $k$. By the Proposition 8 we can always assume that $|V_{k}|=2$. Let us define a coloring of $G$, say $g$, by labelling $v_{i}$ with color $p$ if its corresponding vertex $v_{i}$ is in $V_{p}$. The previous claim ensures that $g$ is a $3$-coloring. Now we show that $g$ is a proper coloring. First note that in $G^{\prime}$, since $e\in V_{p}$, with $p\geq 4$, $e_{a}^{\prime},e_{b}^{\prime}$, and $e_{c}^{\prime}$ must belong to $V_{3},V_{2}$, and $V_{1}$, respectively. Let $e_{t}=v_{i}v_{j}\in E$ and its corresponding vertex $e_{t}$ in $G^{\prime}$ belong to some set $V_{p}$ with $p\geq 4$. This implies that the vertices $\{v_{i},v_{j},v_{e_{t}}\}$ must belong to different sets from $V_{1},V_{2}$, and $V_{3}$. Therefore, $g(v_{i})\neq g(v_{j})$, and hence $g$ is a proper $3$-coloring of $G$. ∎
Hence, we have the following theorem.
Theorem 24.
The Maximum Total Transitivity Decision Problem is NP-complete for bipartite graphs.
5 Polynomial-time algorithm for trees
In the previous section, we have seen that Maximum Total Transitivity Decision Problem is NP-complete for bipartite graphs. In this section, we design a polynomial-time algorithm for finding the total transitivity of a given tree $T=(V,E)$. First, we give a comprehensive description of our proposed algorithm.
5.1 Description of the algorithm
Let $T^{c}$ denote a rooted tree rooted at a vertex $c$, and $T_{v}^{c}$ denote the subtree of $T^{c}$ rooted at a vertex $v$. With a small abuse of notation, we use $T^{c}$ to denote both the rooted tree and the underlying undirected tree, whereas the total transitivity of $T^{c}$ is meant for the underlying undirected tree. To find the total transitivity of $T=(V,E)$, we first define the total transitive number of a vertex $v$ in $T$. The total transitive number of a vertex $v$ in $T$ is the maximum integer $p$ such that $v\in V_{p}$ in a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, where the maximum is taken over all total transitive partitions of $T$. We denote the total transitive number of a vertex $v$ in $T$ by $ttr(v,T)$. Note that if $Tr_{t}(T)=k$ and $v$ is a vertex that belongs to the set $V_{k}$ of a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, then $ttr(v,T)$ must be equal to $k$. Therefore, the total transitivity of $T$ is the maximum total transitive number that a vertex can have; that is, $Tr_{t}(T)=\max\limits_{v\in V}\{ttr(v,T)\}$. Therefore, our goal is to find the total transitive number of every vertex in the tree. Next, we define another parameter, namely the rooted total transitive number. The rooted total transitive number of $v$ in $T^{c}$ is the total transitive number of $v$ in the tree $T_{v}^{c}$, and it is denoted by $ttr^{r}(v,T^{c})$. Therefore, $ttr^{r}(v,T^{c})=ttr(v,T_{v}^{c})$.
Algorithm 1 Total_Transitivity(T)
Input: A tree $T=(V,E)$.
Output: Total transitivity of $T$, that is $Tr_{t}(T)$.
1:Let $V=\{u_{1},u_{2},\ldots,u_{n}\}$ be the set all vertices of $T$;
2:for all $u\in V$ do
3:     Construct a rooted tree $T^{u}$, rooted at a vertex $u$;
4:     Let $\sigma_{u}=(c_{1},c_{2},\ldots,c_{n}=u)$ be the reverse BFS ordering of the vertices of $T^{u}$, rooted at a vertex $u$;
5:     for all $c_{i}$ in $\sigma_{u}$ do
6:         if $c_{i}$ is a leaf then
7:              $mttr^{r}(c_{i},T^{c})=1$;
8:         else
9:              if $c_{i}$ is a support vertex then
10:                  $ttr^{r}(c_{i},T^{c})=mttr^{r}(c_{i},T^{c})=1$;
11:              elseTotal_Transitive_Number$(l_{i_{1}}^{\prime},l_{i_{2}}^{\prime},\ldots,l_{i_{k}}^{\prime})$.        /*where $l_{i_{j}}^{\prime}=l_{i_{j}}$ if $l_{i_{j}}=ttr^{r}(c_{i_{j}},T^{u})$ otherwise $l_{i_{j}}^{\prime}=l_{i_{j}}-1$*/             /* where $c_{i_{1}},\ldots,c_{i_{k}}$ are the children of $c_{i}$ and $l_{i_{1}}^{\prime}\leq l_{i_{2}}^{\prime}\leq\ldots\leq l_{i_{k}}^{\prime}$*/
12:              end if
13:         end if
14:     end for
15:     if We have rooted total transitive number of the root vertex then
16:         $ttr(u,T)=ttr^{r}(u,T^{u})$;
17:     else
18:         $ttr(u,T)=mttr^{r}(u,T^{u})-1$;
19:     end if
20:end for
21:$Tr_{t}(T)$= $\max\limits_{u_{i}\in V}\{ttr(u_{i},T)\}$;
22:return ($Tr_{t}(T)$);
To this end, we define another two parameters, namely the modified-total transitive number and rooted modified-total transitive number. The modified-total transitive number of a vertex $v$ in $T$ is the maximum integer $p$ such that $v\in V_{p}$ in a modified-total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$. Note that the maximum is taken over all modified-total transitive partitions of $T$. We denote the modified-total transitive number of a vertex $v$ in $T$ by $mttr(v,T)$. Finally, we define another parameter, namely the rooted modified-total transitive number. The rooted modified-total transitive number of $v$ in $T^{c}$ is the modified-total transitive number of $v$ in the tree $T_{v}^{c}$, and it is denoted by $mttr^{r}(v,T^{c})$. Therefore, $mttr^{r}(v,T^{c})=mttr(v,T_{v}^{c})$.
Note that the value of the rooted total transitive number of a vertex is dependent on the rooted tree, whereas the total transitive number is independent of the rooted tree. Also, for the root vertex $c$, $ttr^{r}(c,T^{c})=ttr(c,T)$. We recursively compute the rooted total transitive number or rooted modified-total transitive number of the vertices of $T^{c}$ in a bottom-up approach. First, we consider a vertex ordering $\sigma$, the reverse of the BFS ordering of $T^{c}$. For a leaf vertex $c_{i}$, we set $mttr^{r}(c_{i},T^{c})=1$ and for a support vertex $c_{j}$, $ttr^{r}(c_{j},T^{c})=mttr^{r}(c_{j},T^{c})=1$. For other non-leaf vertex $c_{i}$, which is not a support vertex, we call the function Total_Transitive_Number()()( ), which takes the rooted total transitive number or rooted modified-total transitive number of children of $c_{i}$ in $T^{c}$ as input and returns the rooted total transitive number of $c_{i}$ in $T^{c}$ or rooted modified-total transitive number $c_{i}$ in $T^{c}$ according to Theorem 27. At the end of the bottom-up approach, we have the rooted total transitive number of $c$ in $T^{c}$ or rooted modified-total transitive number of $c$ in $T^{c}$, that is, $ttr^{r}(c,T^{c})$ or $mttr^{r}(c,T^{c})$, which is the same as the total transitive number of $c$ in $T$, that is, $ttr(c,T)$ or $ttr(c,T)+1$. After the bottom-up approach, we have the total transitive number of the root vertex $c$ and the rooted total transitive number or rooted modified-total transitive number of every other vertex in $T^{c}$. Similarly, we compute the total transitive number of all vertices of $T$ by considering a rooted tree rooted at vertex $x$. Finally, we compute the maximum overall total transitive number of vertices in $T$ and return that value as the total transitivity of $T$. The algorithm, Total_Transitivity(T), is outlined in Algorithm 1.
5.2 Proof of correctness
In this subsection, we prove the correctness of Algorithm 1. It is clear that the correctness of Algorithm 1 depends on the correctness of the function Total_Transitive_Number()()( ). First, we show the following two lemmas and one theorem, which prove the correctness of the Total_Transitive_Number()()( ) function.
Lemma 25.
Let $v$ be a vertex of $T$, and $ttr(v,T)=t$. Then there exists a total transitive partition of $T$, say $\{V_{1},V_{2},\ldots,V_{i}\}$, such that $v\in V_{i}$, for all $1\leq i\leq t$.
Proof.
Since $ttr(v,T)=t$, there is a total transitive partition $\pi=\{U_{1},U_{2},\ldots,U_{t}\}$ of $T$ such that $v\in U_{t}$. For each $1\leq i\leq t$, let us define another total transitive partition $\pi^{\prime}=\{V_{1},V_{2},\ldots,V_{i}\}$ of $T$ as follows: $V_{j}=U_{j}$ for all $1\leq j\leq(i-1)$ and $V_{i}=\displaystyle{\bigcup_{j=i}^{t}U_{j}}$. Clearly, $\pi^{\prime}$ is a total transitive partition of $T$ of size $i$ such that $v\in V_{i}$. Hence, the lemma follows. ∎
Similar lemma can be defined for modified-total transitivity.
Lemma 26.
Let $v$ be a vertex of $T$, and $mttr(v,T)=t$. Then there exists a modified-total transitive partition of $T$, say $\{V_{1},V_{2},\ldots,V_{i}\}$, such that $v\in V_{i}$, for all $1\leq i\leq t$.
Proof.
Since $mttr(v,T)=t$, there is a modified-transitive partition $\pi=\{U_{1},U_{2},\ldots,U_{t}\}$ of $T$ such that $v\in U_{t}$. Similarly, we can show that there exists a modified-total transitive partition of $T$, say $\{V_{1},V_{2},\ldots,V_{i}\}$, such that $v\in V_{i}$, for all $1\leq i\leq t$. ∎
In the next theorem, we find the total transitive number or modified-total transitive number of a non-leaf and non-support vertex using the total transitive number or modified-total transitive number of its children.
Theorem 27.
Let $v_{1},v_{2},\ldots,v_{k}$ be the children of $x$ in a rooted tree $T^{x}$, and for each $1\leq i\leq k$, $l_{i}$ be either $ttr^{r}(v_{i},T^{x})$ or $mttr^{r}(v_{i},T^{x})$. If $l_{s}$ is both $ttr^{r}(v_{s},T^{x})$ and $mttr^{r}(v_{s},T^{x})$ for some $s$, in that case we consider $l_{s}$ as $ttr^{r}(v_{s},T^{x})$ only. Furthermore, let $l_{j}^{\prime}=l_{j}$, if either $l_{j}=ttr^{r}(v_{j},T^{x})$ otherwise $l_{j}^{\prime}=l_{j}-1$ and $l_{1}^{\prime}\leq l_{2}^{\prime}\leq\ldots\leq l_{k}^{\prime}$. Assume $z$ be the largest integer such that there exists a subsequence of $\{l_{i}^{\prime}:1\leq i\leq k\}$, say, $(l_{i_{1}}^{\prime}\leq l_{i_{2}}^{\prime}\leq\ldots\leq l_{i_{z}}^{\prime})$ such that $l_{i_{p}}^{\prime}\geq p$, for all $1\leq p\leq z$. Also, assume $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Then, either total transitive number or total modified-transitive number of $x$ in the underlying tree $T$ is as follows.
  1. (a)
    If there exists $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such, that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Then we have total transitive number of $x$ in the underlying tree, $T$, and it will be $z+1$, that is $ttr(x,T^{x})=1+z$.
  2. (b)
    If no such $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ exits, such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Then we have modified-total transitive number of $x$ in the underlying tree, $T$, and it will be $z+1$, that is $mttr(x,T^{x})=1+z$.
Proof.
(a) First, consider the case when there exists $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Now, for each $j$, $1\leq j\leq z$, let us consider the subtrees $T_{v_{i_{j}}}^{x}$. It is also given that $ttr^{r}(v_{i_{j}},T^{x})=l_{i_{j}}^{\prime}$, for $j\in\{1,2,\ldots,z\}$. For all $1\leq p\leq z$, since $l_{i_{p}}^{\prime}\geq p$, by Lemma 25, we know that there exists a total transitive partition $\pi^{p}=\{V_{1}^{p},V_{2}^{p},\ldots,V_{p}^{p}\}$ of $T_{v_{i_{p}}}^{x}$, such that $v_{i_{p}}\in V_{p}^{p}$. Since $q$ exits and $mttr^{r}(v_{q},T^{x})=l_{q}=l_{q}^{\prime}+1\geq z+1$, by Lemma 26, we know that there exists a modified-total transitive partition $\pi^{q}=\{V_{1}^{q},V_{2}^{q},\ldots,V_{z+1}^{q}\}$ of $T_{v_{q}}^{x}$, such that $v_{q}\in V_{z+1}^{q}$. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{z},V_{z+1}\}$ of $T^{x}$ as follows: $V_{i}=(\displaystyle{\bigcup_{j=i}^{z}V_{i}^{j}})\cup V_{i}^{q}$, for $2\leq i\leq z$, $V_{z+1}=\{v_{q},x\}$ and every other vertices of $T^{x}$ are put in $V_{1}$. Since $x$ is not a leaf vertex or a support vertex, $\pi$ is a total transitive partition of $T^{x}$. Therefore, by the definition of total transitive number $ttr(x,T^{x})\geq 1+z$.
Next, we show that $ttr(x,T^{x})$ cannot be more than $1+z$. If possible, let $trr(x,T^{x})\geq 2+z$. Then, by Lemma 25, we have that there exists a transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{2+z}\}$ such that $x\in V_{2+z}$. This implies that for each $1\leq i\leq 1+z$, $V_{i}$ contains a neighbour of $x$, say, $v_{i}$ such that the rooted total transitive number of $v_{i}$ is greater or equal to $i$, that is, $l_{i}=ttr^{r}(v_{i},T^{x})\geq i$. The set $\{l_{i}|1\leq i\leq 1+z\}$ forms a desired subsequence of $\{l_{i}:1\leq i\leq k\}$, contradicting the maximality of $z$. Hence, $ttr(x,T)=1+z$.
(b) Let us assume no such $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ exits, such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$. Now, for each $1\leq j\leq z$, let us consider the subtrees $T_{v_{i_{j}}}^{x}$. It is also given that $ttr^{r}(v_{i_{j}},T^{x})=l_{i_{j}}^{\prime}$, for $j\in\{1,2,\ldots,z\}$. For all $1\leq p\leq z$, since $l_{i_{p}}^{\prime}\geq p$, by Lemma 25, we know that there exists a total transitive partition $\pi^{p}=\{V_{1}^{p},V_{2}^{p},\ldots,V_{p}^{p}\}$ of $T_{v_{i_{p}}}^{x}$, such that $v_{i_{p}}\in V_{p}^{p}$. Let us consider the partition of $\pi=\{V_{1},V_{2},\ldots,V_{z},V_{z+1}\}$ of $T^{x}$ as follows: $V_{i}=\displaystyle{\bigcup_{j=i}^{z}V_{i}^{j}}$, for $2\leq i\leq z$, $V_{z+1}=\{x\}$ and every other vertices of $T^{x}$ are put in $V_{1}$. Since $x$ is not a leaf vertex or a support vertex, $\pi$ is a modified-total transitive partition of $T^{x}$. Therefore, $mttr(x,T)\geq 1+z$. Similarly, as above, we can show that $mttr(x,T)=1+z$. ∎
Let $x$ be neither a leaf vertex nor a support vertex of a rooted tree $T^{c}$. Also, assume $v_{1},v_{2},\ldots,v_{k}$ are the children of $x$ in a rooted tree $T^{x}$, and for each $1\leq i\leq k$, $l_{i}$ be either $ttr^{r}(v_{i},T^{x})$ or $mttr^{r}(v_{i},T^{x})$. If $l_{s}$ is both $ttr^{r}(v_{s},T^{x})$ and $mttr^{r}(v_{s},T^{x})$ for some $s$, in that case we consider $l_{s}$ as $ttr^{r}(v_{s},T^{x})$ only. Furthermore, let $l_{j}^{\prime}=l_{j}$, if $l_{j}=ttr^{r}(v_{j},T^{x})$, otherwise, $l_{j}^{\prime}=l_{j}-1$ and $l_{1}^{\prime}\leq l_{2}^{\prime}\leq\ldots\leq l_{k}^{\prime}$. Now from Lemmas 25, 26 and Theorem 27, we have the following algorithm.
Algorithm 2 Total_Transitive_Number$(l_{1}^{\prime},l_{2}^{\prime},\ldots,l_{k}^{\prime})$
Input: $l_{1}^{\prime},l_{2}^{\prime},\ldots,l_{k}^{\prime}$ with $l_{1}^{\prime}\leq l_{2}^{\prime}\leq\ldots\leq l_{k}^{\prime}$.
Output: Either $ttr(x,T^{x})$ or $mttr(x,T^{x})$.
1:Find the largest integer $z$ such that there exists a subsequence of $\{l_{i}^{\prime}:1\leq i\leq k\}$, say, $(l_{i_{1}}^{\prime}\leq l_{i_{2}}^{\prime}\leq\ldots\leq l_{i_{z}}^{\prime})$ such that $l_{i_{p}}^{\prime}\geq p$, for all $1\leq p\leq z$.
2:Find $q\in\{i_{z}+1,i_{z}+2,\ldots,k\}$ such that $l_{q}$ is a $mttr^{r}(v_{q},T^{x})$ and $l_{q}=l_{q}^{\prime}+1\geq z+1$.
3:if $q$ exists then
4:     $ttr(x,T^{x})=z+1$;
5:else
6:     $mttr(x,T^{x})=z+1$;
7:end if
8:return ($mttr(x,T^{x})$ or $ttr(x,T^{x})$ );
5.3 Complexity Analysis
In the function Total_Transitive_Number(), we find the total transitive number or modified-total transitive number of a vertex $x$ based on the rooted total transitive number or rooted modified-total transitive number of its children. We assume the children are sorted according to their $l^{\prime}$-value, which is defined in Theorem 27. Since the for loop in lines $1-8$ of Total_Transitive_Number() runs for every child of $x$, this function takes $O(deg(x))$ time.
In the main algorithm Total_Transitivity(T), the vertex order mentioned in line $4$ can be found in linear time. Then, in a bottom-up approach, we calculate the rooted total transitive numbers or rooted modified-total transitive numbers of every vertex. For that, we are spending $O(deg(c_{i}))$ for every $c_{i}\in\sigma_{u}$. Note that we must pass the children of $c_{i}$ in a sorted order to Total_Transitive_Number(). But as discussed in [14] (an algorithm for finding the Grundy number of a tree), we do not need to sort all the children based on their rooted total transitive numbers or rooted modified-total transitive number; sorting the children whose rooted total transitive number or rooted modified-total transitive number is less than $deg(c_{i})$ is sufficient. We can argue that this can be done in $O(deg(c_{i}))$, as shown in [14]. Hence, the loop in lines $4-11$ takes linear time. To calculate the total transitive number for all vertices, we need $O(n^{2})$ times. Therefore, we have the following theorem:
Theorem 28.
The Maximum Total Transitivity Problem can be solved in polynomial time for trees.
6 Conclusion and future works
In this paper, we have studied the notion of total transitivity in graphs, which is a variation of transitivity. First, we have characterized split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Moreover, for split graph $G$ and $11\leq p\leq\omega(G)-$ 1, we have shown some necessary conditions for $Tr_{t}(G)=p$. Furthermore, we have shown that the decision version of this problem is NP-complete for bipartite graphs. On the positive side, we have proved that this problem can be solved in polynomial time for trees.
This paper concludes by addressing some of the several unresolved problems in the study of total transitivity of a graph.
  1. 1.
    Characterize split graphs with $Tr_{t}(G)=p$ for $p\neq 1,\omega(G)-1$.
  2. 2.
    We know from [18] that in linear time we can solve the transitivity problem in split graphs. Can we design an algorithm for total transitivity in a split graph?
It would be interesting to investigate the complexity status of this problem in other graph classes, such as bipartite chain graphs, bipartite permutation graphs, interval graphs, etc. Designing an approximation algorithm for this problem would be another challenging open problem.